OpenStudy (anonymous):

what is a frequency of an em wave with a wavelength of 30 meters?

OpenStudy (anonymous):

since electromagnetic waves travel at the speed of light, the velocity, v = c = 299,792,458 m/s. the frequency equation is defined as: \[f = \frac{ c }{ \lambda}\] so, \[f = \frac{ 299,792,458 m }{ 30 m*s } = 9,993,081.93 Hz\]